OpenStudy (anonymous):

2(3x-7)+4(3x+2)=6(5x+9)+3

OpenStudy (dumbcow):

distribute and combine like terms

OpenStudy (anonymous):

then move your x's to a side and your constants to the other side

OpenStudy (anonymous):

can you please show me how? I am so confused

OpenStudy (dumbcow):

2(3x+7) multiply the 2 to each term inside the parenthesis 2*3x + 2*7 6x + 14 do that for each one

OpenStudy (anonymous):

so the next one would be 4x3+4x2 = 12x+8

OpenStudy (dumbcow):

good

OpenStudy (anonymous):

the next is 6x5+6x9 = 30x+54

OpenStudy (dumbcow):

good now add terms that have the same variable ex: 2x-4x = -2x you cant add constants with x's

OpenStudy (dumbcow):

i just noticed i messed up on the first one it should be 6x-14

OpenStudy (anonymous):

I saw its okay. I wrote it down as subtraction

OpenStudy (anonymous):

so my problem now shows 6x-14+12x+8=30x+54+3 right

OpenStudy (dumbcow):

correct

OpenStudy (anonymous):

6x+12x+30x = 48x

OpenStudy (anonymous):

If I put the x's together

OpenStudy (dumbcow):

almost dont forget theres an "=" sign separating them so add 6x+12x = 18x but keep 30x on other side for now

OpenStudy (anonymous):

okay so then it becomes......6x+12x=18x+30x+54

OpenStudy (anonymous):

6x-12

OpenStudy (anonymous):

ugh....no its.....18x-14+12=30x+54+3

OpenStudy (dumbcow):

good now add the constants together on both sides

OpenStudy (anonymous):

18x+26=30x+57

OpenStudy (dumbcow):

the +12 should be+8

OpenStudy (dumbcow):

and the 14 is negative

OpenStudy (anonymous):

-14+12= -2

OpenStudy (dumbcow):

correct but it should -14+8 i think you just copied it down wrong

OpenStudy (anonymous):

oops...my problems got mixed up some how.....let me back up. It should be 18x-14+8=30x+54+3

OpenStudy (dumbcow):

correct

OpenStudy (anonymous):

okay so now I get......18x-6=30x+57

OpenStudy (dumbcow):

correct

OpenStudy (dumbcow):

now we want to combine terms from the other side of the equal sign lets add 6 to both sides this works because -6+6=0 on left side

OpenStudy (anonymous):

so left side is just now 18x=30x+57+6 which is 8x=30x+63

OpenStudy (dumbcow):

very good now to get all the x's on the left we should do what?

OpenStudy (anonymous):

would I subtract them since the left side was subtracting??

OpenStudy (dumbcow):

yes, the goal is to cancel the x's on the right

OpenStudy (dumbcow):

30x-30x =0

OpenStudy (anonymous):

18x=63 thats what I am left with right?

OpenStudy (dumbcow):

remember before we added 6 to both sides same here whatever you do to one side you do to the other so we have to subtract 30x on the left too

OpenStudy (anonymous):

12x=63

OpenStudy (dumbcow):

watch out for your signs (18-30=-12) 2/2 = 1 -4/-4 =1 so if we divide a num by itself we always get 1 and the goal is to have 1x on the left side but right now we have -12 so have to divide by -12 on both sides

OpenStudy (anonymous):

x=63/12

OpenStudy (anonymous):

so if we divide 63 by 12 it comes to 5.25. so x=5.25

OpenStudy (dumbcow):

almost watch those signs its negative but yes the number is correct :)

OpenStudy (anonymous):

thank you!

OpenStudy (dumbcow):

your welcome remember those steps